The perimter of a rectangle is 156m. The length is 3m more than twice the width. Find the dimensions.
What is the length ??m
What is the width ??m
Thanks!

Answer by edjones(8007)

L=2w+3
2L+2w=P
2(2w+3)+2w=156
4w+6+2w=156
6w=150
w=25 m
L=2(25)+3
=53 m
.
Ed